superkid2006:

whats the root of 144?

Sailor:

When you say "route" do you mean square roots?

superkid2006:

yes

Sailor:

Oh ok. Do you have to find it by hand or find it using a calculator?

superkid2006:

no calculator

Sailor:

Ok. Can you divide 144 into halves or thirds?

snowflake0531:

if you have to do this by hand, the square root of 144, (hopefully you know what square root means), 10^2 = 100, 11^2 = 121, 12^2=144 so the square root of 144 would be 12, because 12 squared to 144 also, it's good to be able to memorize squares from 1-20*